# HGCN(O): A Self-Tuning GCN HyperModel Toolkit for Outcome Prediction in Event-Sequence Data

### v1

@@ Abstract
journal: Knowledge-Based Systems
@@ 4.1 Node and Graph Attributes Notation and Encoding
[…]
As far as typing is concerned, attributes are categorized into two main types: categorical attributes, encoded using one-hot encoding, and numerical attributes, processed through min-max scaling. When dealing with irrelevant values associated with specific attributes $B_{i}$ that apply only to certain nodes, different strategies are used depending on the attribute type. For numerical attributes, these values are replaced with the median to mitigate the impact of skewness in the data. For categorical attributes, irrelevant values are encoded as 1-1- 1, which also serves as padding. During the training process, padding values are masked to ensure that they do not affect the performance of the model.
@@ 5.5 Hyperparameter Configuration and Optimization Criteria
Key hyperparameters for both convolutional models include the number of GCN and fully connected layers, unit configurations, activation functions, and optional mechanisms like batch normalization and dropout. The architectures also permit modifications in pooling operations, L1 regularization, learning rate schedules, optimizer selection with associated parameters, batch size, and loss functions. For the GraphConv model, we optimize neighborhood aggregation methods due to the absence of an adjacency matrix. In the two TE-GCN hypermodels, embedding dimensions are adjusted. The specific ranges and types of tuned hyperparameters are detailed in Table 1. For all models, we meticulously handled masked values (represented by 1-1- 1) to mitigate biases introduced by irrelevant entries in specific columns of the node vectors.
@@ C O-GCNConv O-GraphConv T-GCNConv T-GraphConv TP-GCNConv TP-GraphConv TE-GCNConv TE-GraphConv S
[…]
@@ 7.2 Models with Balanced Datasets
O-C(A) 1 32 3 211 ReLU F 0.1597 add mean 1 189 ReLU 0.6328 7.977e-3 0.1636 Adam 9.233e-4 6.863e-3 Poly MM 9.753e-5
@@
O-C(O) 1 32 5 111 tanh F 0.5339 9.711e-3 0.3955 max max 1 210 sp 0.3473 2.428e-3 0.3955 Adam 2.223e-3 3.593e-3 Step MM 7.391e-5
@@
T-C(A) 1 32 2 151 tanh F 0.2745 1.149e-3 0.469 mean max 1(S) 111 tanh RMS 7.234e-4 4.415e-3 Cy MM 5.292e-4
@@
T-C(O) 1 64 5 197 ReLU F 0.2854 add add 3(S) 247 ReLU 0.0678 2.667e-3 Adam 9.804e-3 5.032e-3 OCL MM 1.449e-5
@@
TE-C(A) 1 128 4(E) 235 ELU add mean 3(S) 131 ReLU 0.1011 Adam 4.579e-4 2.018e-3 RP MM 3.315e-4
@@ 7.2.2 Hyperparameters
In examining activation functions across GCN layers, we find that Leaky_ReLU and GELU are preferred over softplus and $tanh$, suggesting that their effective gradient flow and non-linearity were optimized by our tuning, enhancing performance in both balanced and imbalanced datasets. This consistency across dataset types underscores our hypermodels’ robustness in supporting stable learning dynamics regardless of class distribution. Embedding models notably apply lower dropout rates in GCN layers, which appears to support generalization while minimizing overfitting. Additionally, minor variations in epsilon and momentum across models suggest these parameters have limited performance impact during tuning, while well-calibrated dropout rates in embedding models further reinforce stability.

### v2

@@ Abstract
journal: Knowledge-Based Systemst1t1footnotetext: Code repository: https://github.com/skyocean/HGCN
@@ 4.1 Node and Graph Attributes Notation and Encoding
Node attributes ($N_{i}$) refer to individual events $X_{i}$ and are categorized into two groups: universal and specific attributes. Universal attributes ($U_{i}$), where $U_{i}\subset N_{i}$, apply to all nodes and are denoted as $U^{a}_{i},U^{b}_{i},U^{c}_{i}$, etc. These attributes are relevant across all nodes, regardless of their neighborhood. In contrast, specific attributes ($B_{i}$), where $B_{i}\subset N_{i}$, apply only to certain nodes and are denoted as $B^{a}_{i},B^{b}_{i},B^{c}_{i}$, etc. The relevance of these attributes is conditional on the values of other attributes. For example, if the value of a universal attribute “process step” is a “register”, the attribute “credit score” is not relevant and would take a placeholder value (“NR”). However, if the value of the “process step” is the “check insurance”, the value of the “credit score” becomes pertinent. Furthermore, since events in sequential logs often have a key attribute such as a unique code or token, we conventionally refer to such attribute as “activity”, and denote it for each node as $\mathcal{A}_{i}$, where $\mathcal{A}_{i}\cap U_{i}=\o$ and $\mathcal{A}_{i}\cap B_{i}=\o$ . The exact name of this attribute will vary depending on the data set or application. For time-stamped event sequences, we also calculate the duration (in seconds) of each “activity”, denoted as $T^{d}_{i}$, where $T^{d}_{i}\subset U_{i}$ and $T^{d}_{i}=T^{c}_{i}-T^{s}_{i}$. Here, $T^{s}_{i}$ represents the start timestamp and $T^{c}_{i}$ represents the end timestamp.
As far as typing is concerned, attributes are categorized into two main types: categorical attributes, encoded using one-hot encoding, and numerical attributes, processed through min-max scaling. When dealing with irrelevant values associated with specific attributes $B_{i}$ that apply only to certain nodes, different strategies are used depending on the attribute type. For numerical attributes, these values are replaced with the median to mitigate the impact of skewness in the data. For categorical attributes, irrelevant values are encoded as 1-1, which also serves as padding. During the training process, padding values are masked to ensure that they do not affect the performance of the model.
@@ 5.5 Hyperparameter Configuration and Optimization Criteria
Key hyperparameters for both convolutional models include the number of GCN and fully connected layers, unit configurations, activation functions, and optional mechanisms like batch normalization and dropout. The architectures also permit modifications in pooling operations, L1 regularization, learning rate schedules, optimizer selection with associated parameters, batch size, and loss functions. For the GraphConv model, we optimize neighborhood aggregation methods due to the absence of an adjacency matrix. In the two TE-GCN hypermodels, embedding dimensions are adjusted. The specific ranges and types of tuned hyperparameters are detailed in Table 1. For all models, we meticulously handled masked values (represented by 1-1) to mitigate biases introduced by irrelevant entries in specific columns of the node vectors.
@@ C O-GCNConv O-GraphConv T-GCNConv T-GraphConv TP-GCNConv TP-GraphConv TE-GCNConv TE-GraphConv S
    Adam: Adam ($\beta_{1}$, $\beta_{2}$); SGD: SGD(momentum); RMS:RMSprop($\alpha$, momentum, eps); Step: Step(step size, $\gamma$); Exp: Exponential($\gamma$); RP: Reduce-on-Plateau(factor, patience, threshold, eps); Poly: Polynomial(total_iters, power); Cos: Cos(eta_min, T_max); OCL: One Cycle(max, pct_start, total_steps); Cy: Cyclic(base, max, step_size_up)
@@ 7.2 Models with Balanced Datasets
O-C(A) 1 32 3 211 ReLU F 0.1597 add mean 1 189 ReLU 0.6328 7.977e-3 0.1636 Adam 9.233e-4 6.863e-3 Poly $M$ 9.753e-5
@@
O-C(O) 1 32 5 111 tanh F 0.5339 9.711e-3 0.3955 max max 1 210 sp 0.3473 2.428e-3 0.3955 Adam 2.223e-3 3.593e-3 Step $M$ 7.391e-5
@@
T-C(A) 1 32 2 151 tanh F 0.2745 1.149e-3 0.469 mean max 1(S) 111 tanh RMS 7.234e-4 4.415e-3 Cy $M$ 5.292e-4
@@
T-C(O) 1 64 5 197 ReLU F 0.2854 add add 3(S) 247 ReLU 0.0678 2.667e-3 Adam 9.804e-3 5.032e-3 OCL $M$ 1.449e-5
@@
TE-C(A) 1 128 4(E) 235 ELU add mean 3(S) 131 ReLU 0.1011 Adam 4.579e-4 2.018e-3 RP $M$ 3.315e-4
@@ 7.2.2 Hyperparameters
In examining activation functions across GCN layers, we find that Leaky_ReLU and GELU are preferred over softplus and tanhtanh, suggesting that their effective gradient flow and non-linearity were optimized by our tuning, enhancing performance in both balanced and imbalanced datasets. This consistency across dataset types underscores our hypermodels’ robustness in supporting stable learning dynamics regardless of class distribution. Embedding models notably apply lower dropout rates in GCN layers, which appears to support generalization while minimizing overfitting. Additionally, minor variations in epsilon and momentum across models suggest these parameters have limited performance impact during tuning, while well-calibrated dropout rates in embedding models further reinforce stability.